OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6. 1 4 5 25

OpenStudy (haseeb96):

3x + 3 = -2x + 2 + 6 3x + 2x = 8 -3 5x = 5 x = 5/5 x = 1 is your answer